La Mia’s Casas builds replicas of residences of famous and infamous people. The company is highly automated, and the new accountant-owner has decided to use machine hours as the basis for predicting maintenance costs. The following data are available from the company’s most recent eight months of operations:

Machine Hours Maintenance Costs
4,000 $1,470
7,000 1,200
3,500 1,680
6,000 1,100
3,000 1,960
9,000 880
8,000 1,020
5,500 1,200

a. Using the high-low method, determine the cost formula for maintenance costs with machine hours as the basis for estimation.

Total cost = $Answer + $Answer X

b. Indicate whether each of the following statements is true or false based on the results of part (a).

(1) Results indicate that as machine hours increase, maintenance cost decreases. TrueFalse
(2) The components of fixed and variable costs are unclear. TrueFalse
(3) Based on the data provided, the equation in part (a) functions correctly. TrueFalse

Tom’s Shoe Repair provides a variety of shoe repair services. Analysis of monthly costs revealed the following cost formulas when direct labor hours are used as the basis of cost determination:

Supplies y = $0 + $4.00X
Production supervision and direct labor y = $500 + $7.00X
Utilities y = $350 + $5.40X
Rent y = $450 + $0.00X
Advertising y = $75 + $0.00X

a. Prepare a flexible budget at 250, 300, 350, and 400 direct labor hours.

250 300 350 400
Variable costs:
Supplies @ $4.00 per DLH
Direct labor @ $7.00 per DLH
Utilities @ $5.40 per DLH
Fixed costs:
Direct labor
Utilities
Rent
Advertising
Total cost

b. Calculate a total cost per direct labor hour at each level of activity. Round amounts to two decimal places.

250 300 350 400
Cost per DLH

c. Tom’s employees usually work 350 direct labor hours per month. The average shoe repair requires 1.25 labor hours to complete. Tom wants to earn a 40 percent margin on his cost. What should be the average charge per customer, rounded to the nearest dollar to achieve Tom’s profit objective?
$Answer
Answer #1

La Mia's Casas

Machine-hours Maintenance Costs
High level 9000 880
Low level 3000 1960
Change 6000 -1080

Variable cost per machine hour = $-1080/6000 = $-0.18

Fixed cost (at high level) = $880 - (9000 x $-0.18) = $880 + $1620 = $2500

Fixed cost (at low level) = $1960 - (3000 x $-0.18) = $1960 + $540 = $2500

a. Total cost = $2500 - $0.18x

b.

(1) Results indicate that as machine hours increase, maintenance cost decreases. TRUE
(2) The components of fixed and variable costs are unclear. FALSE
(3) Based on the data provided, the equation in part (a) functions correctly. TRUE

Per HOMEWORKLIB RULES the first question has been answered. Please post the remaining independent question separately. Thank you.